Jessica Grewal decided to open a food truck business, the Samosa Shack. She encountered the following transactions in managinJournal entry worksheet < 1 2 3 4 5 6 7 8 ... 12 > Record to repairs for truck. Note: Enter debits before credits. Date GenerJournal entry worksheet < 1 2 3 4 5 6 7 8 ..... 12 Record insurance expense of 2015 for 9 months. Note: Enter debits before cJournal entry worksheet < 1 2 3 4 5 6 7 8 ..... 12 Record to depreciation of truck. Note: Enter debits before credits. GeneraJournal entry worksheet < 1 2 3 4 5 6 7 8 ..... 12 > Record depreciation of oven. Note: Enter debits before credits. GeneralJournal entry worksheet < 1 2 3 4 5 6 7 8 ..... 12 > Record to purchase of tires and insurance for year. Note: Enter debits bJournal entry worksheet < 1 2 3 4 5 6 7 8 ..... 12 ) Record to 1 year of insurance expense. Note: Enter debits before creditsJournal entry worksheet < 1 2 3 4 5 6 7 8 ... 12 > Record to depreciation of truck. Note: Enter debits before credits. GeneraJournal entry worksheet < 1 ..... 4 5 6 7 8 9 ... 12 Record to depreciation of oven. Note: Enter debits before credits. GenerJournal entry worksheet < 1 ..... 5 6 7 8 9 10 .. 12 Record to partial year depreciation of Truck. Note: Enter debits beforeJournal entry worksheet < 1 ... 5 6 7 8 9 10 11 12 Record to partial year depreciation of Oven. Note: Enter debits before creJournal entry worksheet < 1 ..... 5 6 7 you to con 8 9 10 11 12 Record to loss on sale of truck. Note: Enter debits before cr
Answer #1

Answer -

1.

Date General Journal Debit Credit
April 01, 2015 Food truck $63000
Oven $6000
Prepaid insurance $6000
   Cash $75000

Calculation:

Food truck = $60000 + $3000 = $63000

2.

Date General Journal Debit Credit
Oct. 01, 2015 Repairs expense $3000
   Cash $3000

3.

Date General Journal Debit Credit
Dec. 31, 2015 Insurance expense $4500
   Prepaid insurance $4500

Calculation:

Prepaid insurance = $6000 * (9/12 months) = $4500

4.

Date General Journal Debit Credit
Dec. 31, 2015 Depreciation expense-Truck $7650
   Accumulated depreciation-Truck $7650

Calculation:

Depreciation expense = (Cost of assets - Residual value) / Useful life

Depreciation expense = ($63000 - $12000) / 5 years

Depreciation expense = $10200 (Annual)

For 9 months = $10200 * (9/12 months) = $7650

5.

Date General Journal Debit Credit
Dec. 31, 2015 Depreciation expense-Oven $765
   Accumulated depreciation-Oven $765

Calculation:

Depreciation expense = (Cost of assets - Residual value) / Useful life

Depreciation expense = ($6000 - $900) / 5 years

Depreciation expense = $1020 (Annual)

For 9 months = $1020 * (9/12 months) = $765

6.

Date General Journal Debit Credit
April. 01, 2016 Prepaid insurance $6000
Repairs expense $3000
   Cash $9000

7.

Date General Journal Debit Credit
Dec. 31, 2016 Insurance expense $6000
   Prepaid insurance $6000

8.

Date General Journal Debit Credit
Dec. 31, 2016 Depreciation expense-Truck $10200
   Accumulated depreciation-Truck $10200

Calculation:

Depreciation expense = (Cost of assets - Residual value) / Useful life

Depreciation expense = ($63000 - $12000) / 5 years

Depreciation expense = $10200

9.

Date General Journal Debit Credit
Dec. 31, 2016 Depreciation expense-Oven $1020
   Accumulated depreciation-Oven $1020

Calculation:

Depreciation expense = (Cost of assets - Residual value) / Useful life

Depreciation expense = ($6000 - $900) / 5 years

Depreciation expense = $1020

10.

Date General Journal Debit Credit
March. 31, 2017 Depreciation expense-Truck $2550
   Accumulated depreciation-Truck $2550

Calculation:

Depreciation expense = (Cost of assets - Residual value) / Useful life

Depreciation expense = ($63000 - $12000) / 5 years

Depreciation expense = $10200 (Annual)

For 3 months = $10200 * (3/12 months) = $2550

11.

Date General Journal Debit Credit
March. 31, 2017 Depreciation expense-Oven $255
   Accumulated depreciation-Oven $255

Calculation:

Depreciation expense = (Cost of assets - Residual value) / Useful life

Depreciation expense = ($6000 - $900) / 5 years

Depreciation expense = $1020 (Annual)

For 3 months = $1020 * (3/12 months) = $255

12.

Date General Journal Debit Credit
March. 31, 2017 Cash $27000
Loss on disposal $19560
Accumulated depreciation-Truck $20400
Accumulated depreciation-Oven $2040
   Food truck $63000
   Oven $6000

Calculation:

Accumulated depreciation-Truck (For 2 years) = Annual depreciation * 2 years

= $10200 * 2 years

= $20400

Accumulated depreciation-Oven (For 2 years) = Annual depreciation * 2 years

= $1020 * 2 years

= $2040

Loss on disposal = [Food truck + Oven] - [Cash + Accumulated depreciation-Truck + Accumulated depreciation-Oven]

= [$63000 + $6000] - [$27000 + $20400 + $2040]

= $19560